OpenStudy (anonymous):

State the reciprocal . -7/35

OpenStudy (shane_b):

-35/7

OpenStudy (shane_b):

Just flip the fraction.

OpenStudy (zzr0ck3r):

yes just flip it and (-a)/b = a/(-b) = -(a/b)